OpenStudy (anonymous):

for times a number is six more than three lines a second number while eight lines the first number is 22 less than seven lines the second number .find the two numbers

OpenStudy (anonymous):

lines=times?

Directrix (directrix):

Let x = the first number and y = the other number. 4x = 3y + 6 8x = 7y - 22

Directrix (directrix):

This is a pair of simultaneous equations: two equations in two variables. Would you multiply the first equation by (-2) and post the result here? Thanks.

Directrix (directrix):

-8x = -6y - 12 8x = 7y -22 ---------------- 0 = y - 34 34 = y ---------> 4x= 3y + 6 4x = 3(34) - 22 4x = 102 + 6 x = 27 ---->